Are the points P = (0,0,−1), Q = (0,1,0) and R = (0,0,1) collinear ?
Answer #1

We have distance formula for A (X1,1,2) and Bltg Mg 1. Zg) - AB= VH2-x)²+(4-7)²4 (29-2,)? now P=(0,0,-1) Q-10,1,0) R= (0,0,1)